OpenStudy (anonymous):

The sides of the cubes are numbered 1 through 6. If they are both tossed, what is the probability that they will both be a 3?

OpenStudy (anonymous):

Assuming you have 2 cubes. P(Both roll a 3) = P( one rolling a 3) * P(2nd one rolling a 3) = 1/6 * 1/6 = 1/36